In 10 years, A will be twice as old as B was 10 years ago. If A is now 9 years older than B, the present age of B is
Answer by mathhelp@(10)

Problem:
In 10 years, A will be twice as old as B was 10 years ago. If A is now 9 years older than B, the present age of B is
lets use A=x
B=y
Now: x = y+9
In ten years: x = 2y-10 = x+10=2y
lets solve!
x = 2y-10
x = y+9
ok, so you have to make one of the equations cancel out x or y
lets cancel out x by multiplying everything by the negative sign (-)
x = 2y-10
-x = -y-9
-----------
0 = y-19
so y is 19
lets find x
2(19)-10 = x
x is 38-10, which is 28